The cost for hosting a dinner party at a restaurant are as follows, write a linear equation to represent the cost, and find the cost for 100 people
people cost
10 300
20 400
50 700

Answer by oberobic(2304)

We can look at these as (x,y) pairs.
(10,300)
(20,400)
(50,700)
.
We can use two of the pairs to define a straight line and then use the third point to force the line go through the third point.
.
m = slope = change in y / change in x
.
300-400 = -100
20-10 = -10
m = 10
.
y = 10x + b
.
700 = 10(50) +b
b = 700-500 = 200
.
y = 10x + 200
.
That means the total cost for the dinner party is $200 + $10 per person.
.
So, x=100 leads us to forecast the total cost = 200 + 10*100 = $1200
.
Done.